# Token Turing Machines are Efficient Vision Models

### v1

@@ Abstract
We propose Vision Token Turing Machines (ViTTM), an efficient, low-latency, memory-augmented Vision Transformer (ViT). Our approach builds on Neural Turing Machines and Token Turing Machines, which were applied to NLP and sequential visual understanding tasks. ViTTMs are designed for non-sequential computer vision tasks such as image classification and segmentation. Our model creates two sets of tokens: process tokens and memory tokens; process tokens pass through encoder blocks and read-write from memory tokens at each encoder block in the network, allowing them to store and retrieve information from memory. By ensuring that there are fewer process tokens than memory tokens, we are able to reduce the inference time of the network while maintaining its accuracy. On ImageNet-1K, the state-of-the-art ViT-B has median latency of 529.5ms and 81.0% accuracy, while our ViTTM-B is 56% faster (234.1ms), with 2.4$\times$ fewer FLOPs, with an accuracy of 82.9%. On ADE20K semantic segmentation, ViT-B achieves 45.65mIoU at 13.8 frame-per-second (FPS) whereas our ViTTM-B model acheives a 45.17 mIoU with 26.8 FPS (+94%).
@@ 1 Introduction
Vision Transformers (ViTs) [8] are used for a variety of vision tasks such as image recognition, semantic segmentation, object detection, and even image generation. ViTs are a building block for many applications and models, as both backbones and foundation models. However, high-performing ViTs incur large computational costs, due to their quadratic complexity with respect to input size. Computational costs can be reduced using token sparsification: the removal of uninformative tokens during inference [27, 1]. However, sparsification compromises the accuracy and, fine-tuning or architectural modifications are required to recover accuracy.
Figure 1: Comparison of our architecture with state-of-the-art methods. It is evident that ViTTM-B(28,28) has lower latency than its accuracy equivalents (e.g. Lookup-ViT7×7) while having higher accuracy than its latency equivalents (e.g. Lookup-ViT3×3).
Architectural components such as memory mechanisms can improve accuracy in various tasks [14, 38, 26]. They are fixed size, and prior work has effectively used memory to enhance models on sequential language and algorithmic tasks [14]. ViTs trained with memory can significantly improve their accuracy [32]. Recently, Token Turing Machines (TTMs) [30] have employed memory with Video Transformers to efficiently process video. LookupViT [20] uses lookup tokens (akin to memory) for efficient process images.
@@ 2 Related Work
In this section we present prior work in ViTs, token-sparsification, and memory-augmented neural networks. Sec. 2.1 presents prior work in ViTs and highlight their differences with ViTTM. We presents prior work in token sparsification in Sec. 2.2 highlight the benefits of ViTTM over sparsification. Prior work in memory-augmented neural networks (MANNs) is presented in Sec. 2.3 and we highlight how we adapt MANNs to non-sequential tasks.
@@ 2.1 Vision Transformers
After transformers achieved success in language processing tasks [37], researchers adapted transformer models for computer vision tasks. ViT [8], which shows that transformers are effective on computer vision tasks. Subsequent works have focused on developing: (1) training regimes that improve ViT task performance, and (2) alternative ViT architectures to achieve better accuracy-latency trade-offs.
@@
Second, a variety of ViT architectures have been proposed that offer different, or strictly better, accuracy-latency trade-offs. These include hierarchical ViTs like SwinTransformer and MViT, which create multi-scale feature maps, and CNN-ViT hybrids that incorporate convolutional biases [22, 11]. Among these are two-stream architectures — our work is most similar to these. Cross-ViT [3], one of the first two-stream ViT designs, processes tokens at different patch sizes in separate branches and then merges them through an efficient cross-attention mechanism. ViT-CoMer [41], utilizes bi-directional interactions with CNNs using a fusion block. Reversible Vision Transformers [23], create a two-stream architecture with the goal of decoupling GPU memory requirements from model depth by reducing activation caching during training. A contemporaneous work, LookupViT [20], processes inputs using two sets of tokens, higher resolution (lookup) tokens and compressed tokens, with information being exchanged between them.
@@ 3.2 ViTTM Architecture
We define the memory and process tokens (streams) as $M\in\mathbb{R}^{T\times d}$, and $P\in\mathbb{R}^{K\times d}$, respectively. After embedding, the tokens are processed using $N$ repeating blocks. We index each block with an index $l$. Each of these ViTTM blocks, performs three operations in sequence: (1) read from memory, (2) processing, and (3) write to memory. The read and write operations return “read tokens” and “write tokens”. We define the read and write tokens as $R\in\mathbb{R}^{K\times d}$ and $W\in\mathbb{R}^{T\times d}$, respectively. To ensure efficiency, we depart from ViT’s approach of processing all tokens and instead apply the processing step only to the process tokens, which are fewer in number than the memory tokens (i.e., $K<<T$). The sequence is formally defined as:
@@
    How tokens will be selected for read/write?
@@ 3.2.2 Read-Write Heads
During the read operation, we set $X_{2}=P^{l-1}$, $X_{1}=M^{l}$, and $V^{\prime}=R^{l}$. Conversely, during the write operation, we set $X_{M}=M^{l}$, $X_{1}=P^{l}$, and $V^{\prime}=W^{l}$. In our evalutation we use $c=d/4$, based on our ablations (Tab. 2).
@@ 4.1 Experimental Setup
Training: Training was conducted on high-performance computing systems equipped with up to 40 NVIDIA A100 GPUs (40GB). We follow ViT training recipes laid out in [33]. Specifically, we pre-train our models on ImageNet-21K [29], and subsequently fine-tune them on ImageNet-1K [6]. For semantic segmentation, we initialize our backbone models with the fine-tuned ImageNet-1K weights.
Measurement: FLOP measurements were performed with the fvcore library [10], and latency was measured using the PyTorch benchmark module [25] with a batch size of 256. Segmentation experiments and benchmarking were performed using the MMSegmentation library [5]. Benchmark and evaluations were performed on both a single 80GB A100 GPU and a 24GB A30 GPU. For reproducibility, our code is open source.
@@ 4.2.2 Read-Write Head and Fusion Choice
We explore combinations of read-write heads and fusion mechanisms outlined in Sec. 3.2. Tab. 1 displays the ablation results for read-write head and fusion mechanisms. We observe that all read-write heads provide comprable accuracy, primarily differing in computational cost. As stated in Sec. 3.2.2, we use Linear Attention [18] due to its efficiency and comparable accuracy to accuracy to Cross Attention.
@@ 4.2.3 Effect of Read-Write Head Latent Dimension
Tab. 2 illustrates the impact of varying the latent embedding dimension $c$. Reducing the latent embedding dimension relative to the embedding dimension, $d$, does not reduce the accuracy but can give substantial latency benefits. As shown in Tab. 2, setting $c=d/4$ reduces latency by 100ms with negligible impact on Top-1 accuracy. Given this result, we choose $c=d/4$ for all evaluations in this work.
@@
Table 2: Ablation over the latent embedding dimension, $c$ for the read-write head on ViTTM-B ($d=768$). Accuracy is computed using the memory tokens rather than the process tokens. We use $\uparrow$/$\downarrow$ notation to indicate “larger/smaller is better”.
@@ 4.2.5 Effect of Memory Stream Non-linearity
ViT/DeiT ViT-B/32 88 4.37 138.3 72.2
@@
Two-Stream CrossViT-B 105 21.22 728.1 82.2
@@
Ours ViTTM-B(28,28) 127 7.08 234.1 82.9
ViTTM-B(32,16) 125 7.10 251.5 80.9
Table 5: Comparison of ViTTM with state-of-the-art methods on image classification (ImageNet-1K). ViTTMs are much faster (lower latency) than state-of-the-art methods while matching their accuracy is most cases. Fig. 1 depicts the data in this table. Latency was measured on a 80GB A100 with batch size 256. Note: LookupViT does not have a public implementation, as such we implement a version following the paper.
4.3 Comparison with other work
4.3.1 Image Classification (ImageNet-1K)
ViTTM offers competitive accuracy-latency trade-offs on ImageNet-1K. Tab. 5 presents our evaluation with existing methods. We achieve lower latency while matching or nearly matching accuracy of other methods, thereby expanding the Pareto frontier (Fig. 1). For example, compared with the similar latency LookupViT3×3, ViTTM-B(28,28) has 5.5% greater accuracy. When compared with the similar accuracy LookupViT7×7, our ViTTM-B(28,28) has a 145.1ms (38%) lower latency.
4.3.2 Semantic Segmentation (ADE20K)
ViTTM also achieves competitive results on ADE20K, as illustrated in Tab. 6. Our ViTTM-B(28,28) matches ViT-B/16 on ADE20K semantic segmentation while improving FPS by 37% (23.8 FPS to 32.5 FPS) on an A100 GPU, and by 94% (13.8 FPS to 26.8 FPS) on an A30 GPU.
Model (M) FPS (A100) $\uparrow$ FPS (A30) $\uparrow$ mIoU $\uparrow$
ViT-B/16 23.8 13.8 45.65
ViTTM-B(28,28) (+37%) 32.5 (+94%) 26.8 45.17
ViTTM-B(32,16) (+38%) 32.8 (+94%) 26.7 43.60
Table 6: Results for semantic segmentaion (ADE20K) using the Segmenter [34] segmentation method. FPS measurements are taken on 80GB NVIDIA A100 and 24GB A30 GPUs.
5 Discussion and Future Work
Two-Stream Architectures: It is evident from both prior work and our experiments that two-stream architectures can be used to create both efficient and accurate vision models, e.g., LookupViT, CrossViT, and ViTTMs. However, two-stream architectures present more design choices than ViTs, leaving much to explore about what constitutes an optimal design. For example, although we ablate many design choices (Sec. 4.2), we do not consider differing embedding dimensions per stream, varying depth, or the effect of self-supervised learning. Previous studies [3] have demonstrated the effectiveness of altering embedding dimensions, yet its interplay with other parameters remains unclear. Changing the embedding dimension has been shown to be effective in prior work [3], but how it is related to other parameters is unknown. Moreover, we also find that two-stream architecture can suffer from training instabilities (Sec. 3.2.4). An important direction for future work is the rigorous study of both architectural choices and the training dynamics of two-stream architectures.
6 Conclusion

### v2

@@ Abstract
We propose Vision Token Turing Machines (ViTTM), an efficient, low-latency, memory-augmented Vision Transformer (ViT). Our approach builds on Neural Turing Machines (NTM) and Token Turing Machines (TTM), which were applied to NLP and sequential visual understanding tasks. ViTTMs are designed for non-sequential computer vision tasks such as image classification and segmentation. Our model creates two sets of tokens: process tokens and memory tokens; process tokens pass through encoder blocks and read-write from memory tokens at each encoder block in the network, allowing them to store and retrieve information from memory. By ensuring that there are fewer process tokens, we are able to reduce the inference time of the network while maintaining its accuracy. On ImageNet-1K, the state-of-the-art ViT-B has median latency of 529.5ms and 81.0% accuracy, while our ViTTM-B is 56% faster (234.1ms), with 2.4$\times$ fewer FLOPs, with an accuracy of 82.9%. On ADE20K semantic segmentation, ViT-B achieves 45.65mIoU at 13.8 frame-per-second (FPS) whereas our ViTTM-B model acheives a 45.17 mIoU with 26.8 FPS (+94%).
@@ 1 Introduction
Vision Transformers (ViTs) [8] are used for a variety of vision tasks such as image recognition, semantic segmentation, object detection, and even image generation. ViTs are a building block for many applications and models, as both backbones and foundation models. However, high-performing ViTs incur large computational costs due to their quadratic complexity with respect to input size. Computational costs can be reduced by processing fewer tokens. Methods such as token sparsification remove uninformative tokens during inference [28, 1]. However, sparsification compromises accuracy and fine-tuning or architectural modifications are required to recover accuracy.
Figure 1: Comparison of our architecture with state-of-the-art methods. It is evident that ViTTM-B(64,64) has lower latency than its accuracy equivalents (e.g. Lookup-ViT7×7) while having higher accuracy than its latency equivalents (e.g. Lookup-ViT3×3).
Architectural components such as memory mechanisms can improve accuracy in various tasks [14, 39, 27]. They are fixed size, and prior work has effectively used memory to enhance models on sequential language and algorithmic tasks [14]. ViTs trained with memory can significantly improve their accuracy [33]. Recently, Token Turing Machines (TTMs) [31] have employed memory with Video Transformers to efficiently process video. LookupViT [21] uses lookup tokens (akin to memory) to efficiently process images.
@@ 2 Related Work
In this section we present prior work in ViTs, token-sparsification, and memory-augmented neural networks. Sec. 2.1 presents prior work in ViTs and highlight their differences with ViTTM. We present prior work in token sparsification in Sec. 2.2 and highlight the benefits of ViTTM over sparsification. Prior work in memory-augmented neural networks (MANNs) is presented in Sec. 2.3 and we highlight how we adapt MANNs to non-sequential tasks.
@@ 2.1 Vision Transformers
After transformers achieved success in language processing tasks [38], researchers adapted transformer models for computer vision tasks. ViT [8] shows that transformers are effective on computer vision tasks. Subsequent works have focused on developing: (1) training regimes that improve ViT task performance, and (2) alternative ViT architectures to achieve better accuracy-latency trade-offs.
@@
Second, a variety of ViT architectures have been proposed that offer different, or strictly better, accuracy-latency trade-offs. These include hierarchical ViTs like SwinTransformer and MViT, which create multi-scale feature maps, and CNN-ViT hybrids that incorporate convolutional biases [23, 11]. Among these are two-stream architectures — our work is most similar to these. Cross-ViT [3] one of the first two-stream ViT designs, processes tokens at different patch sizes in separate branches and then merges them through an efficient cross-attention mechanism. ViT-CoMer [42] utilizes bi-directional interactions with CNNs using a fusion block. Reversible Vision Transformers [24] create a two-stream architecture with the goal of decoupling GPU memory requirements from model depth by reducing activation caching during training. A contemporaneous work, LookupViT [21], processes inputs using two sets of tokens, higher resolution (lookup) tokens and compressed tokens, with information being exchanged between them.
@@ 3.2 ViTTM Architecture
We define the memory and process tokens (streams) as $M\in\mathbb{R}^{T\times d}$, and $P\in\mathbb{R}^{K\times d}$, respectively. After embedding, the tokens are processed using $N$ repeating blocks. We index each block with an index $l$. Each of these ViTTM blocks, performs three operations in sequence: (1) read from memory, (2) processing, and (3) write to memory. The read and write operations return “read tokens” and “write tokens”. We define the read and write tokens as $R\in\mathbb{R}^{K\times d}$ and $W\in\mathbb{R}^{T\times d}$, respectively. To ensure efficiency, we depart from ViT’s approach of processing all tokens and instead apply the processing step only to the process tokens. The sequence is formally defined as:
@@
    How should tokens be selected for read/write?
@@ 3.2.2 Read-Write Heads
During the read operation, we set $X_{2}=P^{l-1}$, $X_{1}=M^{l}$, and $V^{\prime}=R^{l}$. Conversely, during the write operation, we set $X_{2}=M^{l}$, $X_{1}=P^{l}$, and $V^{\prime}=W^{l}$. In our evalutation we use $c=d/4$, based on our ablations (Tab. 2).
@@ 4.1 Experimental Setup
Training: We follow ViT training recipes laid out in [34]. Specifically, we pre-train our models on ImageNet-21K [30], and subsequently fine-tune them on ImageNet-1K [6] at 224px image resolution. For semantic segmentation, we initialize ViTTM backbone models with fine-tuned ImageNet-1K weights. We compare against ViT-B/16 backbones fine-tuned at 224px for fair comparison. Our training recipes are included in the Appendix.
Measurement: FLOP measurements were performed with the fvcore library [10], and latency was measured using the PyTorch benchmark module [26] with a batch size of 256. Segmentation experiments and benchmarking were performed using the MMSegmentation library [5]. Benchmark and evaluations were performed on both a single 80GB A100 GPU and a 24GB A30 GPU.
@@ 4.2.2 Read-Write Head and Fusion Choice
We explore combinations of read-write heads and fusion mechanisms outlined in Sec. 3.2. Tab. 1 displays the ablation results for read-write head and fusion mechanisms. We observe that all read-write heads provide comparable accuracy, primarily differing in computational cost. As stated in Sec. 3.2.2, we use Linear Attention [19] due to its efficiency and comparable accuracy to Cross Attention.
@@ 4.2.3 Effect of Read-Write Head Latent Dimension
Tab. 2 illustrates the impact of varying the latent embedding dimension $c$. Reducing the latent embedding dimension relative to the embedding dimension, $d$, does not reduce accuracy, but can provide substantial latency benefits. As shown in Tab. 2, setting $c=d/4$ reduces latency by 100ms with negligible impact on Top-1 accuracy. Given this result, we choose $c=d/4$ for all evaluations in this work.
@@
Table 2: Ablation over the latent embedding dimension, $c$ for the read-write head on ViTTM-B(49,196) ($d=768$). Accuracy is computed using the memory tokens rather than the process tokens. We use $\uparrow$/$\downarrow$ notation to indicate “larger/smaller is better”. We find latency reductions saturate when $c<d/4$.
@@ 4.2.5 Effect of Memory Stream Non-linearity
ViT/DeiT ViT-B/32 [40] 88 4.37 138.3 72.2
ViT-B/16 [40] 87 16.87 529.5 81.0
DeiT-B/16 [36] 87 16.87 529.7 81.8
Two-Stream CrossViT-B [3] 105 21.22 728.1 82.2
CrossViT-18 [3] 43 9.05 374.1 82.5
CrossViT-18$\dagger$ [3] 44 9.50 378.2 82.8
Rev-ViT-B [24] 86 17.49 556.5 81.8
LookupViT3×3 [21] 90 5.26 230.5 77.9
LookupViT5×5 [21] 90 6.94 297.2 81.6
LookupViT7×7 [21] 90 9.45 379.5 83.0
LookupViT10×10 [21] 90 14.80 563.4 83.9
Ours ViTTM-B(64,64) 127 7.08 234.1 82.9
ViTTM-B(49,196) 125 7.10 251.5 80.9
Table 5: Comparison of ViTTM with state-of-the-art methods on image classification (ImageNet-1K). ViTTMs are much faster (lower latency) than state-of-the-art methods while matching their accuracy is most cases. Fig. 1 depicts the data in this table. Latency was measured on a 80GB A100 with batch size 256. Notes: The ViT baseline model is the 224 resolution fine-tuned model from [8], available from timm [40]. LookupViT does not have a public implementation, as such we implement a version following the paper.
4.3 Comparison with other work
4.3.1 Image Classification (ImageNet-1K)
ViTTM offers competitive accuracy-latency trade-offs on ImageNet-1K. Tab. 5 presents our evaluation with existing methods. We achieve lower latency while matching or nearly matching accuracy of other methods, thereby expanding the Pareto frontier (Fig. 1). For example, compared with the similar latency LookupViT3×3, ViTTM-B(64,64) has 5.5% greater accuracy. When compared with the similar accuracy LookupViT7×7, our ViTTM-B(64,64) has a 145.1ms (38%) lower latency.
4.3.2 Semantic Segmentation (ADE20K)
ViTTM also achieves competitive results on ADE20K across multiple devices, as illustrated in Tab. 6. Our ViTTM-B(64,64) matches ViT-B/16224 on ADE20K semantic segmentation while improving FPS by 37% (23.8 FPS to 32.5 FPS) on an A100 GPU, and by 94% (13.8 FPS to 26.8 FPS) on an A30 GPU. In this work, we do not investigate the reason for the discrepancy between devices.
Model (M) FPS (A100) $\uparrow$ FPS (A30) $\uparrow$ mIoU $\uparrow$
ViT-B/16384 [35] 23.8 13.8 48.06
ViT-B/16224 23.8 13.8 45.65
ViTTM-B(64,64) (+37%) 32.5 (+94%) 26.8 45.17
ViTTM-B(49,196) (+38%) 32.8 (+94%) 26.7 43.60
Table 6: Results for semantic segmentation (ADE20K) using the Segmenter [35] segmentation method. FPS measurements are taken on 80GB NVIDIA A100 and 24GB A30 GPUs.
5 Discussion and Future Work
Two-Stream Architectures: It is evident from both prior work and our experiments that two-stream architectures can be used to create both efficient and accurate vision models, e.g., LookupViT, CrossViT, and ViTTMs. However, two-stream architectures present more design choices than ViTs, leaving much to explore about what constitutes an optimal design. For example, although we ablate many design choices (Sec. 4.2), we do not consider differing embedding dimensions per stream, varying depth, or the effect of self-supervised learning. Previous studies [3] have demonstrated the effectiveness of altering embedding dimensions, yet its interplay with other parameters remains unclear. Changing the embedding dimension has been shown to be effective in prior work [3], but how it is related to other parameters is unknown. Moreover, we also find that two-stream architectures can suffer from training instabilities (Sec. 3.2.4). An important direction for future work is the rigorous study of both architectural choices and the training dynamics of two-stream architectures.
6 Conclusion
Appendix
In this appendix, we provide details of our training regime and provide additional evaluation on ImageNet1K. First, in Sec. 6.1, we detail hyperparameters and other details about our training regime for ViTTM. Second, in Sec. 6.2, we compare ViTTMs with Token Merging [1], and show results for a small ViTTM model (ViTTM-S). Our code can be found at https://github.com/pjjajal/EfficientTTMs.
6.1 Training Configurations
In Sec. 4.1, we briefly described our training recipe for ViTTM. Here, Tab. 7 provides details of both our pre-training and fine-tuning regimes to enhance reproducibility of our work. ”RRC” indicates the use of random resize crop. The ”CE” Loss refers to Cross Entropy, and ”BCE” refers to Binary Cross Entropy.
Pre-training Fine-Tuning
Eff. Batch size 4096 2048
Optimizer AdamW AdamW
LR $1.5\times 10^{-4}$ $0.25\times 10^{-4}$
Warmup LR $1.0\times 10^{-6}$ $1.0\times 10^{-6}$
Min. LR $1.0\times 10^{-7}$ $1.0\times 10^{-7}$
LR decay cosine cosine
Weight decay 0.03 0.1
Warmup epochs 3 5
Epochs 300 300
Stoch. Depth $0.1$ $0.1$
Gradient Clip. $1.0$ $1.0$
Image Size 224 224
Horiz. flip $\checkmark$ $\checkmark$
RRC $\checkmark$ $\checkmark$
RandAug Ops $\times$ $2$
RandAug Mag. $\times$ $20$
Mixup alpha $\times$ $0.8$
CutMix alpha $\times$ $0.8$
Erasing prob. $\times$ $0.25$
Loss CE BCE
Table 7: Training configurations for ViTTM-B models. All training was performed on NVIDIA A100 80GB GPU’s.
6.2 Extra Results
ViT/DeiT ViT-S/16 22 4.25 149.5 74.2
DeiT-S/16 22 4.25 152.0 79.8
ViT-B/32 88 4.37 138.3 72.2
@@ 6.2 Extra Results
Two-Stream CrossViT-S 27 5.63 235.7 81.0
CrossViT-15 28 5.81 249.1 82.3
CrossViT-15$\dagger$ 28 6.13 252.3 81.5
CrossViT-B 105 21.22 728.1 82.2
@@
Token Merging [1] ViT-S/16(r=2) 22 4.31 172.7 74.0
ViT-S/16(r=4) 22 4.02 161.6 73.8
ViT-S/16(r=8) 22 3.41 138.2 73.1
ViT-S/16(r=10) 22 3.14 125.9 72.5
ViT-S/16(r=12) 22 2.85 115.2 71.6
ViT-S/16(r=14) 22 2.57 103.3 70.4
ViT-S/16(r=16) 22 2.30 94.0 68.1
ViT-B/16(r=2) 86 16.46 551.1 81.0
ViT-B/16(r=4) 86 15.34 515.7 80.9
ViT-B/16(r=8) 86 13.12 440.6 80.4
ViT-B/16(r=10) 86 12.02 402.0 80.1
ViT-B/16(r=12) 86 10.93 367.0 79.6
ViT-B/16(r=14) 86 9.84 330.2 78.9
ViT-B/16(r=16) 86 8.78 296.4 77.6
Ours ViTTM-S(64,64) 33 1.84 77.7 79.2
ViTTM-B(64,64) 127 7.08 234.1 82.9
ViTTM-B(49,196) 125 7.10 251.5 80.9
Table 8: Comparison of ViTTM with state-of-the-art methods on image classification (ImageNet-1K). Latency was measured on a 80GB A100 with batch size 256. Notes: The ViT baseline model is the 224 resolution fine-tuned model from [8], available from timm [40]. LookupViT does not have a public implementation, as such we implement a version following the paper. Token Merging [1] is applied to ViT-S/16 and ViT-B/16 models at various pruning rates ($r$) without fine-tuning.
We present extra ImageNet-1k results in Tab. 8. Specifically, we trained a ViTTM-S model, and include comparisons against ViT-S and ViT-B augmented with Token Merging at various pruning rates (without fine-tunign). ViTTMs consistently have lower latency than state-of-the-art methods while matching their accuracy. Compared with Token Merging, ViTTMs achieve higher accuracy (as expected), while having lower latency across a range of pruning ratios ($r$).